5) Suppose a price floor on sparkling wine is proposed by the Health Minister of the country of Vinyardia. What will be the likely effect (relative to original equilibrium) on the market for sparkling wine in Vinyardia?

a) Quantity demanded will decrease, quantity supplied will increase, and a surplus will result.

b) Quantity demanded will increase, quantity supplied will decrease, and a surplus will result.

c) Quantity demanded will decrease, quantity supplied will increase, and a shortage will result.

d) Quantity demanded will increase, quantity supplied will decrease, and a shortage will result.

6) If the government implements a price ceiling on insulin, this will

a) increase the price consumers will pay for insulin.

b) decrease the quantity of insulin the manufacturers will be willing to supply.

c) have to be set above the market equilibrium price to be effective.

d) encourage manufacturers to produce and sell more of insulin to increase their profits.

7) To affect the market outcome, a binding price ceiling

a) must be set below the black market price.

b) must be set below the legal price.

c) must be set below the price floor.

d) must be set below the equilibrium price.

8) The demand and supply of honey are represented by the following equations where P is the price in dollars per unit of honey and Q is the quantity of honey:

Market Demand for honey:    Q= 20– (1/2)P

Market Supply of honey:        Q= P - 4

Given the above information and holding everything else constant, what are the values of Consumer Surplus (CS) and Producer Surplus (PS) in the market for honey?

a) CS = $72 and PS = $144

b) CS = $288 and PS = $144

c) CS = $144 and PS = $72

d) CS = $144 and PS = $288
Answer #1

5) Solution: Quantity demanded will decrease, quantity supplied will increase, and a surplus will result

Explanation: A price floor on sparkling wine will reduce the quantity demanded and increase the quantity supplied

 

6) Solution: decrease the quantity of insulin the manufacturers will be willing to supply

Explanation: When the government implements a price ceiling on insulin, it would reduce the insulin's quantity that the manufacturers will be willing to supply

 

7) Solution: must be set below the equilibrium price

Explanation: The price ceiling below equilibrium or price floor above equilibrium is binding

8) Solution: CS = $144 and PS = $72

Working: Market Demand for honey: Q= 20– (1/2)P

Q=20-0.5P

0.5P = 20-Q

P = 40 - 2Q

Market Supply of honey: Q= P - 4 or P = 4+Q

At equilibrium:

20 - 0.5 P = P - 4

P + 0.5P = 20+ 4

1.5P = 24

P = 16

Q = P - 4 = 16 - 4 = 12

Consumer surplus: P at intercept of demand is 40. Thus the height will be 40 - 16 = 24

CS = 1/2* 24 * 12 = 144

Similarly: P at intercept of demand is 4. Thus the height will be 12

CS = 1/2* 12 * 12 = 72